Seventy-five percent of female prison inmates are mothers. If 3 female prison inmates are selected at random, what is the probability that none are mothers? Round the final answer to three decimal places, if necessary.


p(none are mothers)=





1
Expert's answer
2020-07-23T17:46:51-0400

Given : p(mother ) = 0.75

Formula : p(Ac)= 1-p(A)

Solution :

P(non mother )=1-p(mother)

=1-0.75

=0.25


Hence p( three are not mothers)= 0.25*0.25*0.25

= 0.0156= 0.016.